### v1

@@ 5 Analysis
In this section, we analyze the influence of several issues on the performance. Firstly, in Section 3.2.1, several strategies have been introduced to select pixel pairs from the representation $\mathbf{M}^{so}$. The influence of these selection strategies is analyzed. Secondly, in Section 3.2.2, the global consolidation component comprises a channel attention module and a spatial attention module. They can be implemented in a series framework or parallel framework. Finally, a visualization analysis is given to show more details about our model.
@@ 5.1 Influence of Local Consolidation Architecture
The local consolidation module is stacked with several convolution blocks. A convolution block comprises a pixel difference convolution and a normalization layer. Each pixel convolution block can employ a pixel selection strategy. Therefore, the architecture of the local consolidation module can be determined by two elements: the pixel selection strategy and the number of convolution block.
In this experiment, 9 pixel selection strategies (or PDC) is referred as: ’CPDC-XY’ ( ’$C_{xy}$’ in short), ’CPDC-DG’ (’$C_{d}$’), ’CPDC-OMNI’ (’C’ ), ’APDC-CW’ (’$A$’), ‘CNN-2D” ( ’$V$’), ’APDC-CCW’ (’$A_{r}$’), ’RPDC-XY’ (’$R_{xy}$’), ’RPDC-DG’ (’$R_{d}$’), ’RPDC-OMNI’ (’$R$’). The number of convolution blocks ranges from 1 to 4. Under this setting, the architecture of the local consolidation can be represented as a set. For example, $[C_{xy}-C_{d}-A_{r}-R]$ indicates that it is a local consolidation composed of 4 convolution blocks stacked as ’CPDC-XY’ (’$C_{xy}$’), ’CPDC-DG’ (’$C_{d}$’), ’APDC-CCW’ (’$A_{r}$’) and ’RPDC-OMNI’ (’$R$’), respectively. In this experiment, the $\mathrm{WebNLG}^{*}$ dataset is adopted. The Table 6 presents the performance of various architectures, where the top-4 configurations are highlighted in bold.
@@ 3 Convolution Blocks
Architecture $[C-A_{r}-R]$ $\mathbf{[C-A-A_{r}]}$ $[C-C-A_{r}]$
$Prec./Rec./F1$ $93.2/95.4/94.3$ $\mathbf{93.9}/\mathbf{95.5}/\mathbf{94.7}$ $93.1/95.2/94.2$
@@ 4 Convolution Blocks
Architecture $\mathbf{[C_{xy}-C_{d}]\times 2}$ $[A-A_{r}]\times 2$ $[R_{xy}-R_{d}]\times 2$
$Prec./Rec./F1$ $\mathbf{94.0}/\mathbf{95.6}/\mathbf{94.8}$ $93.5/94.9/94.2$ $93.6/95.4/94.5$
@@
Prec./Rec./F1 $93.7/95.2/94.5$ $93.3/95.6/94.5$ $93.2/95.4/94.3$
Architecture $[C_{xy}-C_{d}-A-A_{r}]$ $[C_{xy}-C_{d}-R_{xy}-R_{d}]$ $[A-A_{r}-R_{xy}-R_{d}]$
Prec./Rec./F1 $94.0/95.1/94.6$ $93.8/95.0/94.4$ $94.1/95.2/94.6$
Architecture $[C_{xy}-C_{d}-A_{r}-R]$ $[C-C-A_{r}-R]$ $[C-A-A_{r}-R]$
Prec./Rec./F1 $93.6/95.3/94.5$ $93.8/95.1/94.5$ $94.4/94.8/94.6$
Architecture $[C-A_{r}-R_{xy}-R_{d}]$ $\mathbf{[C-A_{r}-R-R]}$ $[C-A_{r}-A_{r}-R]$
Prec./Rec./F1 $93.9/95.2/94.6$ $\mathbf{93.7}/\mathbf{95.6}/\mathbf{94.7}$ $93.9/95.2/94.6$
Architecture $\mathbf{[C-A_{r}-R-V]}$ $[C_{xy}-C_{d}-R_{d}-V]$ $[C_{xy}-C_{d}-A_{r}-V]$
Prec./Rec./F1 $\mathbf{94.1}$/$\mathbf{96.0}$/$\mathbf{95.0}$ $93.7/95.4/94.5$ $93.8/95.1/94.5$
The result shows that increasing the number of convolution blocks improves the performance. Stacking identical types of PDCs provides additional performance gains without degrading the model. The reason is that multi-stacked convolution blocks have two advantages to support the task. First, multi-stacked convolutions possess the ability to learn higher-level abstract features, which are beneficial in supporting the task of relation triple extraction. Second, stacked blocks can enlarge the receipt field, making it effective for learning semantic dependencies within a sentence.
The Top-4 configurations are ‘$[C-A-A_{r}]$’, ‘$[C_{xy}-C_{d}]\times 2$’, ‘$[C-A_{r}-R-R]$’ and ‘$[C-A_{r}-R-V]$’, which utilize various types of convolutions for capturing semantic information in different encoding directions. Specifically, the stacking of three layers of ‘$[C-A-A_{r}]$’ demonstrates the superiority of bidirectional APDCs, which may leverage semantic contextual information from adjacent elements to aid in entity boundary recognition. Similarly, the utilization of four layers of ‘$[C_{xy}-C_{d}]\times 2$’ not only underscores its prioritization of extracting features from central and adjacent positions but also showcases the substantiated efficiency of acquiring information from both the XY axis and diagonal directions. The characteristic of ‘$[C-A_{r}-R-V]$’ is the stack of four omnidirectional PDCs, which exhibits the highest performance. This can be attributed to the comprehensive utilization of four distinct encoding directions of PDCs. Therefore, the ‘$[C-A_{r}-R-V]$’ configuration is employed as the default for our model.
@@ 5.2 Influence of Global Consolidation Architecture
The global consolidation component contains a channel attention module and a spatial attention module, which can be set in a series or parallel framework. To show the influence of the global consolidation architecture, in this experiment, we evaluate the model’s performance by organizing them in a series or parallel framework. In the series framework, the output of the sentence encoding module is sequentially processed by the channel attention module and spatial attention module, respectively. In the parallel framework, the channel attention and spatial attention are paralleled. Furthermore, in the whole bi-consolidating module, the local consolidation component and the global consolidation component can be placed into a series or a parallel framework too. It is also evaluated in this section. This experiment employs the $\mathrm{WebNLG^{*}}$ dataset. The influence of the architectures is listed in Table 7, where the symbol “&” signifies a parallel connection, while “$+$” is employed to denote a series connection.
@@ 5.3 Visualization Of Different Bi-consolidating Module Components
A 2D sentence representation unfolds a two-dimensional semantic space. This approach effectively expresses the semantic structures of a sentence and provides additional semantic information. The 2D representation is fed into the bi-consolidating module to generate a fine-grained subject-object features, encoded with local and global semantic features. In addition to quantitative analysis, it is also informative to provide a visualized analysis to show the influence of the bi-consolidating module. In this visualization, we collect five sentences with different overlapping patterns as SEO, SOO, Normal, EPO + SOO, SEP + SOO. These sentences are collected from the $\mathrm{WebNLG^{*}}$ dataset. They are introduced as follows:
@@
1) SEO: the sentence is “Agra airport has the location identifier VIAG and is located in Awadh, Uttar Pradesh, Bundelkhand”. It contains two relational triples as: [“Pradesh”, isPartOf, “Awadh”], [“Pradesh”, isPartOf, “Bundelkhand”]. The entity pairs (“Pradesh”, “Awadh”) and (“Pradesh”, “Bundelkhand”), with the overlapping subject “Pradesh”, belong to a single entity overlap.
@@
2) SOO: the sentence is “A Fortress of Grey Ice is from the United States, led by President Barack Obama”. It contains two relational triple as: [“States”, leaderName, “Obama”], [“Ice”, country, “States”]. The entity pairs (“States”, “Obama”) and (“Ice”, “States”) have the entity "State" as their subject and object, respectively, resulting in a subject object overlap.
@@
3) Normal: the sentence is “Elliot See, born July 23 1927 in Dallas, has died”. It contains a single relational triple as: [“See”, birthPlace, “Dallas”]. There is no overlap among them.
@@
4) EPO + SOO: the sentence is “Ampara Hospital is in Eastern Province, Sri Lanka. Ranil Wickremesinghe is a leader there”. It contains three relational triples as: [“Hospital”, country, “Lanka”], [“Lanka”, leaderName, “Wickremesinghe”] and [“Hospital”, state, “Lanka”]. The entity pairs (“Hospital”, “Lanka”) and (“Lanka”, “Wickremesinghe”) have the entity "Lanka" as their subject and object, resulting in a subject object overlap.
@@
5) SEP + SOO: the sentence is “Alan Shepard was a crew member of NASA operated Apollo 14 who died in California which is represented by Dianne Feinstein”. It contains four relational triples as: [“Shepard”, was a crew member of, “14”], [“Shepard”, deathPlace, “California”], [“California”, senators, “Feinstein”] and [“14”, operator, “NASA”]. The entity pairs (“Shepard”, “14”) and (“Shepard”, “California”) belong to a single entity overlap. The entity pairs (“Shepard”, “California”) and (“California”, “Feinstein”) have the entity "California" as their subject and object, resulting in a subject object overlap.
@@
These sentences are first fed into the sentence encoding module to generate $\mathbf{M}^{so}\in\mathbb{R}^{N\times N\times D_{h}}$. Then, the $\mathbf{M}^{so}$ is fed into the bi-consolidating module to learn the task relevant subject-object feature representation $\mathbf{T}^{so}$. In our model, the bi-consolidating module is composed of a local consolidation component and a global consolidation component. The local consolidation contains 4 convolution blocks referred as: “CPDC-OMNI”, "APDC-CCW", "RPDC-OMNI" and "CNN-2D". To show the visualization of various components in the bi-consolidating module, we transform each 2D representation an image by implementing the Global Average Pooling operation. This resulting image can be seen as a heat map, which denotes to a semantic plane. The visualization of these semantic planes in the five sentences is shown in Figure 5.
@@
Figure 5 is divided into five parts marked with “SEO”, “SOO”, “Normal”, “EPO + SOO” and “SEP + SOO”. Each part is composed of five images belonging to a sentence, which are further categorized into local consolidation section and global consolidation section. The local consolidation section shows the visualization of each block as "CPDC-OMNI", "APDC-CCW", "RPDC-OMNI", "CNN-2D".
@@
It is interesting to observe that after the utilization of the convolution “CPDC-OMNI”, all entities exhibit distinct contour characteristics. For instance, in the “Normal”, the entities “See” and “Dallas” have a high intensity. Their span edges exhibit significant semantic intensity changes in the semantic plane. Next, following the convolution block "APDC-CCW", the semantic representation surrounding the entity is amplified, which guarantees the retention of all the semantic information within the sentence. Next to the convolution block "RPDC-OMNI", the semantic information within a neighborhood is intensified, which mitigates any noise that might exist in adjacent elements. Subsequently, after the convolution block "CNN-2D", the contextual semantic representation in the proximity of the entity undergoes further enhancement, simultaneously minimizing potential disturbances originating from neighboring elements.
@@
As visualized in every part, it can be seen that all entities exhibit obvious intensity. On the other hand, non-entity regions manifest comparatively lower intensity. In the visualization of the global consolidation module, the most impressive phenomenon is that all entity semantic representations have clear contour characteristics and have a high intensity in a 2D sentence representation. This also proves that this module has the advantage to learn remote semantic dependencies in a sentence. Another interesting phenomenon is that the representation, composed of the semantic information about an entity pair, also has a concretized semantic representation.

### v2

@@ 5 Analysis
In this section, we analyze the influence of several issues on the performance. Firstly, in Section 3.2.1, several strategies have been introduced to select pixel pairs from the representation $\mathbf{M}^{so}$. The influence of these selection strategies is analyzed. Secondly, in Section 3.2.2, the global consolidation component comprises a channel attention module and a spatial attention module. They can be implemented in a series framework or parallel framework. Thirdly, a deeper analysis of the computational efficiency of the model. Finally, a visualization analysis is given to show more details about our model.
@@ 5.1 Influence of Local Consolidation Architecture
In Figure 4, eight pixel difference convolution operations were presented: CPDC-XY ($C_{xy}$ in short), CPDC-DG ($C_{d}$), CPDC-OMNI ($C$), APDC-CW ($A$), APDC-CCW ($A_{r}$), RPDC-XY ($R_{xy}$), RPDC-DG ($R_{d}$), RPDC-OMNI ($R$). In these operations, $C_{xy}$ and $C_{d}$ are sub-operations of C. Except the dependency direction, A and $A_{r}$ have similar semantic structures. $R_{xy}$ and $R_{d}$ are also sub-operations of R. After adding the normal convolution operation CNN-2D ($V$), these convolutions can be divided into four groups: [$C_{xy}$, $C_{d}$, $C$], [$A$, $A_{r}$], [$R_{xy}$, $R_{d}$, $R$] and [$V$].
A convolution block comprises a pixel difference convolution layer and a normalization layer. The local consolidation component can be stacked with different convolution blocks. Therefore, the architecture of the local consolidation module can be determined by the type of pixel convolutions and the number of blocks. This experiment is conducted to analyse the influence of local consolidation architecture. The performance is evaluated on the $\mathrm{WebNLG}^{*}$ dataset. Table 6 presents the performance of various architectures, where the top-6 configurations are highlighted in bold.
@@ 3 Convolution Blocks
Architecture $[C-A_{r}-R]$ ${[C-A-A_{r}]}$ $[C-C-A_{r}]$
P/R/F1 $93.2/95.4/94.3$ $\mathbf{{93.9}/{95.5}/{94.7}}$ $93.1/95.2/94.2$
@@ 4 Convolution Blocks
Architecture ${[C_{xy}-C_{d}]\times 2}$ $[A-A_{r}]\times 2$ $[R_{xy}-R_{d}]\times 2$
P/R/F1 $\mathbf{{94.0}/{95.6}/{94.8}}$ $93.5/94.9/94.2$ $93.6/95.4/94.5$
@@
P/R/F1 $93.7/95.2/94.5$ $93.3/95.6/94.5$ $93.2/95.4/94.3$
Architecture $[C_{xy}-C_{d}-A-A_{r}]$ $[C_{xy}-C_{d}-R_{xy}-R_{d}]$ $[C_{xy}-C_{d}-R_{d}-V]$
P/R/F1 $94.0/95.1/94.6$ $93.8/95.0/94.4$ $93.7/95.4/94.5$
Architecture $[C_{xy}-C_{d}-A_{r}-R]$ $[C_{xy}-C_{d}-A_{r}-V]$ $[C_{xy}-A_{r}-R-V]$
P/R/F1 $93.6/95.3/94.5$ $93.8/95.1/94.5$ $93.8/95.3/94.5$
Architecture $[C-A_{r}-R_{xy}-V]$ $[A-A_{r}-R_{xy}-R_{d}]$ $[C-A-R-V]$
P/R/F1 $93.6/95.5/94.5$ $94.1/95.2/94.6$ $93.9/95.3/94.6$
Architecture $[C-A_{r}-R_{xy}-R_{d}]$ ${[C-A_{r}-R-R]}$ $[C_{d}-A_{r}-R-V]$
P/R/F1 $93.9/95.2/94.6$ $\mathbf{{93.7}/{95.6}/{94.7}}$ $\mathbf{93.8}/\mathbf{95.8}/\mathbf{94.8}$
Architecture $[C-A-A_{r}-R]$ $[C-A_{r}-R_{d}-V]$ ${[C-A_{r}-R-V]}$
P/R/F1 $94.4/94.8/94.6$ $\mathbf{93.7/95.7/94.7}$ $\mathbf{94.1}$/$\mathbf{96.0}$/$\mathbf{95.0}$
The results in Table 6 show three fundamental principles to build a local consolidation module.
First, increasing the number of convolution block is helpful to improve the performance. Even stacking identical types of PDCs can provide additional performance gains. The reason is that multi-stacked convolution blocks possess the ability to learn higher-level abstract features, which are beneficial in supporting the task of relation triple extraction. Furthermore, stacked blocks can enlarge the receipt field of the input, making it effective for learning semantic dependencies within a sentence.
Second, stacking convolution blocks with different pixel difference operations can achieve better performance. For example, the two configurations $[C_{d}-A_{r}-R-V]$ and ${[C-A_{r}-R-V]}$ have the best performance, which utilize various types of convolutions for capturing semantic information in different encoding directions. But there are exceptions. For example, the bidirectional APDCs $[C-A-A_{r}]$ also demonstrates the superiority of local consolidations, which may leverage semantic contextual information from adjacent elements to aid in entity boundary recognition. $[C_{xy}-C_{d}]\times 2$ not only underscores its prioritization of extracting features from central and adjacent positions, but also showcases the substantiated efficiency of acquiring information from both the XY axis and diagonal directions.
Third, replacing pixel difference convolutions in the same group is less influential on the final performance, e.g., [$C_{xy}$, $C_{d}$, $C$], [$A$, $A_{r}$], [$R_{xy}$, $R_{d}$, $R$]. The reason may be that, in the same group, convolution operations capture semantic structures in the same pattern. Furthermore, in the same group, the OMNI-version operations (e.g., $C$ and $R$) usually have better performance, because other pixel difference convolutions in the same group are usually sub-operations of the OMNI-version operation.
@@ 5.2 Influence of Global Consolidation Architecture
The global consolidation component contains a channel attention module and a spatial attention module, which can be set in a series or parallel framework. To show the influence of the global consolidation architecture, in this experiment, we evaluate the model’s performance by organizing them in a series or parallel framework. In the series framework, the output of the sentence encoding module is sequentially processed by the channel attention module and spatial attention module, respectively. In the parallel framework, the channel attention and spatial attention are paralleled. Furthermore, in the whole bi-consolidating module, the local consolidation component and the global consolidation component can be placed into a series or a parallel framework too. The influence is also evaluated in this section.
In this experiment, the $\mathrm{WebNLG^{*}}$ dataset is employed to evaluate the performance. The influence of the architectures is listed in Table 7, where the symbol “&” signifies a parallel connection, while “$+$” is employed to denote a series connection.
@@
5.3 Computational Efficiency
The bi-consolidating model is mainly composed of a global consolidation component and a local consolidation component. The global consolidation component contains a channel attention module and a spatial attention module, which are normal operations in natural language processing. The local consolidation component is stacked of four convolution blocks, where different pixel difference convolutions are used for learning high-order abstract features. In order to evaluate the computational efficiency of the bi-consolidating model, we compare the training times between five architectures with different local consolidation components: $[C-A_{r}-R-V]$, $[V-V-V-V]$, $[V-A_{r}-R-V]$, $[C-V-R-V]$ and $[C-A_{r}-V-V]$.
In this experiment, for making a fair comparison, instead of directly compared with a deep architecture composed of a convolution layer, the $[V-V-V-V]$ model is used as the baseline, which is composed of four normal convolution layers. This experiment is conducted on the $\mathrm{WebNLG^{*}}$ dataset. The batch size for training and testing are set to 6 and 1, respectively. The max length of input sentence is 100. The result is shown in Figure 5, where the training time (second) means the time required to implement one epoch in the training process.
Figure 5: Computational efficiency.
In Figure 5, the four architectures achieve similar performance. This result indicates that different pixel convolution operations are less influential on the computational complexity. Compared the performance between them, $[V-V-V-V]$ achieves the lowest computational complexity. The reason is that $[V-V-V-V]$ only implements convolution operation. It is effective to support parallel computing. On the other hand, as shown in Figure 4, the $C$, $R$, $A_{r}$ and $A$ convolution operations need to compute the semantic dependencies of elements relevant to adjacent elements, which limit the utilization of parallel computing. Especially, the $R$ operation has a larger kernel size, which further degenerates the final performance. Even so, compared $[C-A_{r}-V-V]$ with $[V-V-V-V]$, it only decreases about 5% computing speed.
@@ 5.4 Visualization Of Different Bi-consolidating Module Components
A 2D sentence representation unfolds a two-dimensional semantic space. This approach effectively expresses the semantic structures of a sentence and provides additional semantic information. The 2D representation is fed into the bi-consolidating module to generate a fine-grained subject-object features, encoded with local and global semantic features. In addition to quantitative analysis, it is also informative to provide a visualized analysis to show the influence of the bi-consolidating module. In this visualization, we collect five sentences with different overlapping patterns as SEO, SOO, Normal, EPO + SOO, SEO + SOO. These sentences are collected from the $\mathrm{WebNLG^{*}}$ dataset. They are introduced as follows:
@@
1) SEO: the sentence is “Agra airport has the location identifier VIAG and is located in Awadh, Uttar Pradesh, Bundelkhand”. It contains two relational triples as: [“Pradesh”, “isPartOf”, “Awadh”], [“Pradesh”, “isPartOf”, “Bundelkhand”]. The entity pairs (“Pradesh”, “Awadh”) and (“Pradesh”, “Bundelkhand”), with the overlapping subject “Pradesh”, belong to a single entity overlap.
@@
2) SOO: the sentence is “A Fortress of Grey Ice is from the United States, led by President Barack Obama”. It contains two relational triple as: [“States”, “leaderName”, “Obama”], [“Ice”, “country”, “States”]. The entity pairs (“States”, “Obama”) and (“Ice”, “States”) have the entity “States” as their subject and object, respectively, resulting in a subject object overlap.
@@
3) Normal: the sentence is “Elliot See, born July 23 1927 in Dallas, has died”. It contains a single relational triple as: [“See”, “birthPlace”, “Dallas”]. There is no overlap among them.
@@
4) EPO + SOO: the sentence is “Ampara Hospital is in Eastern Province, Sri Lanka. Ranil Wickremesinghe is a leader there”. It contains three relational triples as: [“Hospital”, “country”, “Lanka”], [“Lanka”, “leaderName”, “Wickremesinghe”] and [“Hospital”, “state”, “Lanka”]. The entity pairs (“Hospital”, “Lanka”) and (“Lanka”, “Wickremesinghe”) have the entity “Lanka” as their subject and object, resulting in a subject object overlap.
@@
5) SEO + SOO: the sentence is “Alan Shepard was a crew member of NASA operated Apollo 14 who died in California which is represented by Dianne Feinstein”. It contains four relational triples as: [“Shepard”, “was a crew member of”, “14”], [“Shepard”, “deathPlace”, “California”], [“California”, senators, “Feinstein”] and [“14”, “operator”, “NASA”]. The entity pairs (“Shepard”, “14”) and (“Shepard”, “California”) belong to a single entity overlap. The entity pairs (“Shepard”, “California”) and (“California”, “Feinstein”) have the entity “California” as their subject and object, resulting in a subject object overlap.
@@
These sentences are first fed into the sentence encoding module to generate $\mathbf{M}^{so}\in\mathbb{R}^{N\times N\times D_{h}}$. Then, the $\mathbf{M}^{so}$ is fed into the bi-consolidating module to encode the subject and object features of all possible entity pairs in a sentence. In our model, the bi-consolidating module is composed of a local consolidation component and a global consolidation component. The local consolidation contains 4 convolution blocks referred as: “CPDC-OMNI”, “APDC-CCW”, “RPDC-OMNI” and “CNN-2D”. To show the visualization of various components in the bi-consolidating module, we transform each 2D representation an image by implementing the Global Average Pooling operation. This resulting image can be seen as a heat map, which denotes to a semantic plane. The visualization of these semantic planes in the five sentences is shown in Figure 6.
@@
Figure 6 is divided into five parts marked with “SEO”, “SOO”, “Normal”, “EPO + SOO” and “SEO + SOO”. Each part is composed of five images belonging to a sentence, which are further categorized into local consolidation section and global consolidation section. The local consolidation section shows the visualization of each block as “CPDC-OMNI”, “APDC-CCW”, “RPDC-OMNI”, “CNN-2D”.
@@
It is interesting to observe that after the utilization of the convolution “CPDC-OMNI”, all entities exhibit distinct contour characteristics. For instance, in the “Normal”, the entities “See” and “Dallas” have a high intensity. Their span edges exhibit significant semantic intensity changes in the semantic plane. Next, following the convolution block “APDC-CCW”, the semantic representation surrounding the entity is amplified, which guarantees the retention of all the semantic information within the sentence. Next to the convolution block “RPDC-OMNI”, the semantic information within a neighborhood is intensified, which mitigates any noise that might exist in adjacent elements. Subsequently, after the convolution block “CNN-2D”, the contextual semantic representation in the proximity of the entity undergoes further enhancement, simultaneously minimizing potential disturbances originating from neighboring elements.
@@
As visualized in every part, it can be seen that all entities exhibit greater intensity. On the other hand, non-entity regions manifest comparatively lower intensity. In the visualization of the global consolidation module, the most impressive phenomenon is that all entity semantic representations have clear contour characteristics and have a high intensity in a 2D sentence representation. This also proves that this module has the advantage to learn remote semantic dependencies in a sentence. Another interesting phenomenon is that representations, composed of the semantic information about an entity pair, also have concretized semantic representations.